# Randomly punctured Reed–Solomon codes achieve list-decoding capacity over linear-sized fields

### v5

@@ A.1 Preliminaries
A.1.0.0.1 Dual of Reed–Solomon codes.
@@
A.1.0.0.2 Generic Zero Patterns.
Following [BGM23], we leverage the GM-MDS theorem to establish list-decodability of Reed–Solomon codes. In this work, we more directly connect the list-decoding problem to the GM-MDS theorem using a hypergraph orientation lemma (introduced in the next section). Here, we review generic zero-patterns and the GM-MDS theorem. To keep the meaning of the variable “$k$” consistent throughout the paper, we unconventionally state the definition of zero patterns and the GM-MDS theorem with $n-k$ rows instead of $k$ rows.
@@
Given positive integers $k\leq n$, an $(n,n-k)$-generic-zero-pattern (GZP) is a collection of sets $S_{1},\ldots,S_{n-k}\subset[n]$ such that, for all $K\subseteq[n-k]$,
@@
A.1.0.0.3 GM-MDS Theorem.
As in [BGM23], we connect the list-decoding problem to the GM-MDS theorem. Here, we make the connection more directly.
Theorem A.2 (GM-MDS Theorem [DSY14, Lov18, YH19]).
Given $q\geq 2n-k-1$ and any generic zero-pattern $S_{1},\dots,S_{n-k}\subset[n]$, there exists pairwise distinct evaluation points $\alpha_{1},\dots,\alpha_{n}\in\mathbb{F}_{q}$ and an invertible matrix $M\in\mathbb{F}_{q}^{(n-k)\times(n-k)}$ such that, if $H$ is the parity-check matrix for $\mathsf{RS}_{n,k}(\alpha_{1},\dots,\alpha_{n})$ (as in (38)), then $MH$ achieves zero-pattern $S_{1},\dots,S_{n-k}$, meaning that $(MH)_{\ell,i}=0$ whenever $i\in S_{\ell}$.
@@ A.2 Hypergraph orientations
Our new perspective of the tools from [BGM23] leverages a well-known theorem about orienting weakly-partition-connected hypergraphs, stated below. This theorem is most explicitly stated in [Fra11], but it implicit in [Kir03, FKK03a].
@@
Theorem A.3 (Theorems 9.4.13 and 15.4.4 of [Fra11]).
A hypergraph $\mathcal{H}$ is $k$-weakly-partition-connected if and only if it has an orientation such that, for some vertex $v$ (the “root”), every other vertex $u$ has $k$ edge-disjoint paths to $v$ [^9] In [Fra11, Theorems 9.4.13 and 15.4.4], the property of having $k$ edge-disjoint paths to $v$ is called $(0,k)$-edge-connected.
We note that Theorem 9 is remains true if “to $v$” is replaced with “from $v$” and $k$-weakly-partition-connected is replaced with another hypergraph notion called $k$-partition-connected. The following corollary essentially captures (the hard direction of) [BGM23, Lemma 2.8].
@@
Let $\mathcal{H}=([t],\mathcal{E})$ be a $k$-weakly-partition-connected hypergraph with $n$ hyperedges and $t\geq 2$. Then there exists integers $\delta_{1},\dots,\delta_{t}\geq 0$ summing to $n-k$ such that taking $\delta_{j}$ copies of $S_{j}\stackrel{{\scriptstyle\rm def}}{{=}}\{i\in[n]:j\notin e_{i}\}\subset[n]$ gives an $(n,n-k)$-GZP.
@@
Take the orientation of $\mathcal{H}$ and root vertex $v\in[t]$ given by Theorem 9. We now take our $\delta_{j}$’s as follows: for each non-root $j\in[t]$, let $\delta_{j}\stackrel{{\scriptstyle\rm def}}{{=}}\deg_{in}(j)$ to be the in-degree of vertex $j$. For the root $v$, let $\delta_{v}\stackrel{{\scriptstyle\rm def}}{{=}}\deg_{in}(v)-k$. Note that any other vertex $u$ has $k$ edge-disjoint paths to $v$, so $v$ has in-degree at least $k$ and $\delta_{v}\geq 0$. Since there are $n$ hyperedges, the sum of all $\delta_{j}$’s is thus $n-k$. We now check the generic zero pattern condition (39). Consider any nonempty multiset $K\subset[t]$ such that each vertex $j\in[t]$ appears at most $\delta_{j}$ times. We claim:
@@
The first equality holds by definition of $S_{j}$. The second equality holds because $\sum_{j\in[t]}\delta_{j}=n-k$. The second inequality holds because $|K|\leq\sum_{j\in K}\delta_{j}$ by definition of $K$. It reamins to show the first inequality. We have two cases:
Case 1: the root $v$ is in $K$. The number of hyperedges induced by the vertices $[t]\setminus K$ is at most the sum of the indegrees of $[t]\setminus K$, which is exactly $\sum_{j\in[t]\setminus K}\delta_{j}$ by definition of $\delta_{j}$.
Case 2: the root $v$ is in $[t]\setminus K$. Fix an arbitrary vertex $u$ in $K$. By our orientation of $\mathcal{H}$, vertex $u$ has $k$ edge-disjoint paths to $v$. Each of these paths has an edges that “enters” $[t]\setminus K$, i.e., the head is in $[t]\setminus K$ but the edge is not induced by $[t]\setminus K$. Thus, the number of edges induced by $[t]\setminus K$ is at most $(\sum_{j\in[t]\setminus K}\deg_{in}(j))-k$, which is exactly $\sum_{j\in[t]\setminus K}\delta_{j}$ by definition of $\delta_{j}$. Hence, we have the first inequality. This covers all cases, proving (40), completing the proof. ∎
@@
Appendix B Alphabet size limitations
In this section, we establish Proposition 1.5. For positive integers $m$, view $\mathbb{F}_{2^{m}}$ as a vector space of dimension $m$ over base field $\mathbb{F}_{2}$. For a set $S\subset\mathbb{F}_{2^{m}}$, let
Lemma B.1 (Proposition 3.2 of [BKR10]).
Let $L$ be a subspace of $\mathbb{F}_{2^{m}}$. Then $P_{L}$ has the form
$X^{2^{\dim L}}+\sum_{i=0}^{\dim L-1}\alpha_{i}X^{2^{i}}.$ (49)
where $\alpha_{i}\in\mathbb{F}_{2^{m}}$
As an immediate corollary, we have
Lemma B.2.
Let $L$ be an affine subspace of $\mathbb{F}^{2^{m}}$. Then $P_{L}$ has the form
$X^{2^{\dim L}}+\sum_{i=0}^{\dim L-1}\alpha_{i}X^{2^{i}}+\beta$ (50)
for $\alpha_{i},\beta\in\mathbb{F}_{2^{m}}$.
Proof.
Since $L$ is an affine subspace, there exists $\gamma$ such that $L-\gamma\stackrel{{\scriptstyle\rm def}}{{=}}\{\alpha-\gamma:\alpha\in L\}$ is a subspace of $\mathbb{F}_{2^{m}}$. By Lemma B.1, we have $P_{L-\gamma}$ is of the form
$X^{2^{\dim L}}+\sum_{i=0}^{\dim L-1}\alpha_{i}X^{2^{i}}$ (51)
Lemma B.3 (Analogous to Lemma 3.5 of [BKR10]).
Proof.
so there are at least $2^{v(m-v)}$ affine-subspaces $L$ with $|L\cap S|\geq n/2^{m-v}$. For all such affine-subspaces $L$, the polynomial $P_{L}(X)$ has the form $X^{2^{v}}+\sum_{i=0}^{v-1}\alpha_{i}X^{2^{i}}+\beta$ by Lemma B.2. Among these affine-subspaces $L$, by the pigeonhole principle, for at least a fraction $2^{-m(v-u-1)}$ of these subspaces, their subspace polynomials $P_{L}(X)$ have the same $\alpha_{i}$ for $i=u+1,u+2,\dots,v$. Let $\mathcal{L}$ be this family of subspaces. The number of subspaces is at least $2^{v(m-v)}\times 2^{-m(v-u-1)}=2^{(u+1)m-v^{2}}$, so $\mathcal{L}$ is the desired family of affine subspaces. ∎
Proof of Proposition 1.5.
Let $\delta=2^{-r-1}$ as in the statement of Proposition 1.5. Consider a Reed–Solomon code of length $n$ and rate $\delta$ over $\mathbb{F}_{q}$, where $q=2^{m}$ with $m$ sufficiently large. Let $S\subset\mathbb{F}_{q}$ be the set of $n$ evaluation points. Apple Lemma B.3 with $u=m-\lceil{1.99r}\rceil$ and $v=m-r$. This gives a family $\mathcal{L}$ of $2^{m(m-\lceil{1.99r}\rceil)-(m-r)^{2}}=2^{2rm-\lceil{1.99r}\rceil m+r^{2}}\geq q^{\Omega(\log(1/\delta))}$ affine subspaces $L\leq\mathbb{F}_{2^{m}}$ for which $|L\cap S|\geq n/2^{m-v}=2\delta n$. Furthermore, for $L\in\mathcal{L}$, the subspace polynomials $P_{L}$ each have $2^{v}$ roots, and agree on all coefficients of degree larger than $2^{u}$. Let $L_{0}$ be an arbitrary element of $\mathcal{L}$. Then the polynomials $\{P_{L_{0}}-P_{L}:L\in\mathcal{L}\}$ are each of degree at most $2^{u}=2^{-\lceil{1.99r}\rceil}q\leq 4\delta^{1.99}q\leq\delta n$, and each agree with $P_{L_{0}}(X)$ on at least $|L\cap S|\geq 2\delta n$ values of $S$. Thus, our Reed–Solomon code is not $(1-2\delta,n^{\Omega(1/\delta)})$-list-decodable, as desired. ∎

### v6

@@ A.1 Preliminaries
Dual of Reed–Solomon codes.
@@
Generic zero patterns.
Following [9], we leverage the GM-MDS theorem to establish the list-decodability of Reed–Solomon codes. In this work, we more directly connect the list-decoding problem to the GM-MDS theorem using a hypergraph orientation lemma (introduced in the next section). Here, we review generic zero patterns and the GM-MDS theorem. To keep the meaning of the variable “kk” consistent throughout the paper, we unconventionally state the definition of zero patterns and the GM-MDS theorem with nkn-k rows instead of $k$ rows.
@@
Given positive integers $k\leq n$, an (n,nk)(n,n-k)-generic zero pattern (GZP) is a collection of sets $S_{1},\ldots,S_{n-k}\subseteq[n]$ such that, for all $K\subseteq[n-k]$,
@@
GM-MDS Theorem.
As in [9], we connect the list-decoding problem to the GM-MDS theorem. Here, we make the connection more directly.
Theorem A.2 (GM-MDS Theorem [13, 49, 65]).
Given $q\geq 2n-k-1$ and any generic zero pattern $S_{1},\dots,S_{n-k}\subseteq[n]$, there exists pairwise distinct evaluation points $\alpha_{1},\dots,\alpha_{n}\in\mathbb{F}_{q}$ and an invertible matrix $M\in\mathbb{F}_{q}^{(n-k)\times(n-k)}$ such that, if $H$ is the parity-check matrix for $\mathsf{RS}_{n,k}(\alpha_{1},\dots,\alpha_{n})$ (as in (30)), then $MH$ achieves zero pattern $S_{1},\dots,S_{n-k}$, meaning that $(MH)_{\ell,i}=0$ whenever $i\in S_{\ell}$.
@@ A.2 Hypergraph Orientations
Our new perspective of the tools from [9] leverages a well-known theorem about orienting weakly-partition-connected hypergraphs, stated below. This theorem is most explicitly stated in [20], but it is implicit in [45, 22].
@@
Theorem A.3 (Theorems 9.4.13 and 15.4.4 of [20]).
A hypergraph $\mathcal{H}$ is $k$-weakly-partition-connected if and only if it has an orientation such that, for some vertex $v$ (the “root”), every other vertex $u$ has $k$ edge-disjoint paths to vv. [^10] In [20, Theorems 9.4.13 and 15.4.4], the property of having $k$ edge-disjoint paths to $v$ is called $(0,k)$-edge-connected.
We note that Theorem 10 remains true if “to vv” is replaced with “from vv” and $k$-weakly-partition-connected is replaced with another hypergraph notion called $k$-partition-connected. The following corollary essentially captures (the hard direction of) [9, Lemma 2.8].
@@
Let $\mathcal{H}=([t],\mathcal{E})$ be a $k$-weakly-partition-connected hypergraph with $n$ hyperedges and $t\geq 2$. Then there exists integers $\delta_{1},\dots,\delta_{t}\geq 0$ summing to nkn-k such that taking $\delta_{j}$ copies of $S_{j}\stackrel{{\scriptstyle\rm def}}{{=}}\{i\in[n]:j\notin e_{i}\}\subseteq[n]$ gives an (n,nk)(n,n-k)-GZP.
@@
Take the orientation of $\mathcal{H}$ and root vertex $v\in[t]$ given by Theorem 10. We now take our $\delta_{j}$’s as follows: for each non-root $j\in[t]$, let $\delta_{j}\stackrel{{\scriptstyle\rm def}}{{=}}\deg_{in}(j)$ to be the in-degree of vertex $j$. For the root $v$, let $\delta_{v}\stackrel{{\scriptstyle\rm def}}{{=}}\deg_{in}(v)-k$. Note that any other vertex $u$ has $k$ edge-disjoint paths to $v$, so $v$ has in-degree at least $k$ and $\delta_{v}\geq 0$. Since there are $n$ hyperedges, the sum of all $\delta_{j}$’s is thus nkn-k. We now check the generic zero pattern condition (31). Consider any nonempty multiset $K\subseteq[t]$ such that each vertex $j\in[t]$ appears at most $\delta_{j}$ times. We claim:
@@
The first equality holds by definition of $S_{j}$. The second equality holds because $\sum_{j\in[t]}\delta_{j}=n-k$. The second inequality holds because $|K|\leq\sum_{j\in K}\delta_{j}$ by definition of $K$. It remains to show the first inequality. We have two cases:
Case 1: the root $v$ is in $K$. The number of hyperedges induced by the vertices $t]\setminus K$ is at most the sum of the indegrees of $t]\setminus K$, which is exactly $\sum_{j\in[t]\setminus K}\delta_{j}$ by definition of $\delta_{j}$.
Case 2: the root $v$ is in $t]\setminus K$. Fix an arbitrary vertex $u$ in $K$. By our orientation of $\mathcal{H}$, vertex $u$ has $k$ edge-disjoint paths to $v$. Each of these paths has an edge that “enters” $t]\setminus K$, i.e., the head is in $t]\setminus K$ but the edge is not induced by $t]\setminus K$. Thus, the number of edges induced by $t]\setminus K$ is at most $(\sum_{j\in[t]\setminus K}\deg_{in}(j))-k$, which is exactly $\sum_{j\in[t]\setminus K}\delta_{j}$ by definition of $\delta_{j}$. Hence, we have the first inequality. This covers all cases, proving (32), completing the proof. ∎
@@
Acknowledgments
We thank Mary Wootters and Francisco Pernice for helpful discussions about [9] and the hypergraph perspective of the list-decoding problem. We thank Karthik Chandrasekaran for helpful discussions about hypergraph connectivity notions and for the reference of Theorem 10 in [20]. We thank Nikhil Shagrithaya and Jonathan Mosheiff for pointing out a mistake in the proof of Lemma 4.3 in an earlier version of the paper. We thank an anonymous reviewer for pointing out a mistake in the proof of Corollary A.4 in an earlier version of the paper.